Parth (parthkohli):

If 100 coins are tossed, what is the probability that exactly 50 heads will be showing?

Parth (parthkohli):

Is this as simple as saying "1/2"?

Parth (parthkohli):

Since I found this from the 50 Hard Probability Problems webpage( http://www.delphiforfun.org/Programs/Math_Topics/FiftyProbabilityProblems.htm), I think there is something tricky going on there.

OpenStudy (mimi_x3):

Binomial Probability?

Parth (parthkohli):

I am not sure.

OpenStudy (mimi_x3):

100C50 * (1/2)^{50} * (1/2)^{50}

Parth (parthkohli):

Hmm, but could you possibly explain that a little?

OpenStudy (mimi_x3):

\[\tbinom{100}{50} *\left(\frac{1}{2}\right)^{50}*\left(\frac{1}{2}\right)^{50} = 0.0796\] \(p=q= \frac{1}{2}\) This is a particular case of probability of binomial probability when probabilities \(p\) and \(q\) of successes and failure are both \(\frac{1}{2}\)

OpenStudy (anonymous):

You have to get 50 heads and 50 tails the probability of that is \[ \frac 1 {2^{50}}\frac 1 {2^{50}} \] There are \[ 100C50 \] such possiblities

Parth (parthkohli):

Oh, yes! How could I forget combinations?!

OpenStudy (anonymous):

The answer is about .08

OpenStudy (mimi_x3):

woops..but am i wrong?

Parth (parthkohli):

Now I really feel silly.

Parth (parthkohli):

Thank you @eliassaab @Mimi_x3 Mimi, I might not be a qualified person to judge that. :p

OpenStudy (anonymous):

yw

OpenStudy (mimi_x3):

@eliassaab: I have a question..I don't think I'm wrong..this looks like Binomial Probability..

OpenStudy (mimi_x3):

wait forget it i read the question wrongly..100 coins are tossed..not 100 coins are tossed 100 times..sorry.

OpenStudy (anonymous):

Mimi, you are right.

OpenStudy (mimi_x3):

I'm sorry; but I'm confused I'm not that good with Probability. Re-reading the question again it says "100 coins are tossed" does that mean that it's not Binomial Probability? Since it is not 100 times?

OpenStudy (anonymous):

You can think about one coin tossed 100 times.

OpenStudy (mimi_x3):

But 100 coins are tossed; does that mean it is all tossed at the same time? So it is not Binomial Probability.

OpenStudy (anonymous):

Tossing 100 coins and tossing a coin 100 times is the same event.

OpenStudy (mimi_x3):

ok. i got it. thanks.